ASSET Cash 50.000.000 Current assets 200.000.000 Total Assets 250.000.000 Liabilities Short term debt 55.000.000 Current Liabilities 55.000.000 Long term debt 120.000.000 Preffered Stock 12.000.000 Common Stock 18.000.000 Retairned Earnings 24.000.000 Total Common Equity 36.000.000 Liabilities and Equity 250.000.000 short term debt is 2.000.000 EUR bank loan with 496 interest.During this year is close to zero Long term debt consists of 8% bonds with market price 82 Euro and Face value 100 Eur.YTM is 12% Preferred stock has par value of 100 Euros, divident yield 12% and trade price (market price) 92 Euros Company has 14.000.000 common shares.Market price is 20 Eur.Last divident payment was 2.400.000 Euro . · . and divident are expected to grow with rate 8% stock beta is 1.5 and the yield of bond is 7% , while market risk (Rm-Rf) is 6% Tax rate is 30% . · Calculate Wacc of the company based on book values and market values
Answer #1

WACC = (E x Ke) + ((Kd x (1-T) x D) + (P x Kp)

E + D + P

Where:

E = market value / Book Value of the firm’s equity (common stock + retained earning)
D = market value/ / Book Value of the firm’s debt
P = market value / Book Value of the firm’s Preference Stock
Kp = cost of Preference Stock
Ke = cost of equity (required rate of return)
Kd = cost of debt (yield to maturity on existing debt)
T = tax rate

But

Kp = D/P0

Where:

P0 = market value of preference stock
D = Preference dividend

So Kp = 12/92 =13.04 %

but in case of WACC on book value value P0 shall be Book value & so Kp shall be 12/100 =12 %

&

Kd = I x (1-T)

Where:

T = tax rate
I = Interest Received on debt

So Kd (Long term debt) = 8 x (1-0.3) =5.60 %

but in case of WACC on market value interest on debt shall be replaced by YTM so Kd shall be 12 x (1-0.3) =8.4 %

For Ke

Ke = (D1 / P0 ) + g

Where:

P0 = market /book value of common Equity (Retained Earning & common stock)
D1 = Dividend of next year i.e Dividend of current year x (1+ g)

g = Growth rate of dividend

So Ke (Book value) = 2400000 x (1+0.08) / 14000000 + 0.08 = 15.20 %

36000000/14000000

Ke (Market value) = 2400000 x (1+0.08) / 14000000 + 0.08 = 8.93 %

20

So WACC (book value) =     (E x Ke) + ((Kd x (1-T) x D) + (P x Kp)

E + D + P

= (36,000,000 x 15.20%) + [(5.60 % x (1-0.3)) x 120,000,000] + (12% x 12,000,000)   = 6.914 %

36,000,000 +120,000,000 +12,000,000

& WACC (market value) =     (E x Ke) + ((Kd x (1-T) x D) + (P x Kp)

E + D + P

= ((14,000,000 x 20) x 8.93%) + [(8.40 % x (1-0.3)) x (1,200,000 x 82)] + (13.04% x (120,000 x 92))   = 8.276%

(14,000,000 x 20) + (1,200,000 x 82) +(120,000 x 92)